How do we find a line parallel to y=4x-5 and passes through (-2,3)? I'm helping my child with homework. 4 is the slope, right? I'm not sure what the next step is.
Answer by Alan3354(69443)

How do we find a line parallel to y=4x-5 and passes through (-2,3)? I'm helping my child with homework. 4 is the slope, right? I'm not sure what the next step is.
===========================
4 is the slope.
m = 4
-----
y-y1 = m*(x - x1) where (x1,y1) is the point.
y-3 = 4*(x+2) is an equation.
y = 4x + 11 also